The price of a large pizza decreased from $21.00 to $17.00. As a result, the quantity demanded of skateboards increased from 300.00 to 320.00.

1st attempt

Part 1   (1 point)

Feedback

See Hint

Using the midpoint formula, what is the percentage change observed for the price of a large pizza? Give your answer to two decimal places.
----------%

Part 2   (1 point)

Feedback

See Hint

Using the midpoint formula, what is the percentage change observed for the quantity demanded of skateboards? Give your answer to two decimal places.
-------- %

Part 3   (1 point)

See Hint

What is the cross-price elasticity of demand between pizzas and skateboards? Give your answer to two decimal places.
Answer #1

Part 1.

Percentage change in price = (P2-P1) /[(P2+P1) /2]

=(17-21)/[(17+21)/2]

=-4/19

=0.2105

=-21.05%

Part 2

Percentage change in quantity =(Q2-Q1) /[(Q2+Q1) /2]

=(320-300)/[(300+320)/2]

=20/310

=0.645

=6.45%

Part - 3

Cross price elasticity of demand =percentage change in quantity of skateboard /percentage change in price of large pizza.

=-21.05/6.45

=-3.26